CutQua company expects to incur $800,000 in overhead costs this coming year- $200,000 in thee Cut and Polish department and S600,000 in the Quality Control department. Total annual direct labor costs are expected to be $160,000. The Cut and Polish department expects to use 20,000 machine hours, and the Quality Control department plans to utilize 40,000 hours of direct labor time for the year. Total available hours (capacity) are 25,000 for the Cut and Polish artment and 60,000 for the Quality Control department.

Required: 1) Assume CutQua Company allocates overhead costs with the plant-wide approach, and direct labor cost is the allocation base. Calculate the rate used by the company to allocate overhead costs.

2) Assume CutQua Company allocates overhead costs with the department approach using alternately budgeted and capacity amounts for the allocation bases. Calculate the rate used by each department to allocate overhead costs.

3) Determine the total amount of unused capacity costs implied if the predetermined rate is computed using the capacity figures.

4)Determine the actual cost rates if at the end of the period: a. Overhead costs incurred are $210,000 in the Cut and Polish department and $590,000 in the Quality Control department. b. Direct labor costs are S 150,000. c. Machine hours used by Cut and Polish department are 21,000. d. Direct labor hours utilized within the Quality Control department are 48,000.
Answer #1
1) Overhead allocation rate = $800000/$160000 = 500.00% of labor cost
2) CUT & POLISH DEPARTMENT:
Allocation rate based on capcaity = $200000/25000MH = $              8.00 per MH
Allocation rate based on budgeted activity = $200000/20000MH = $            10.00 per MH
QUALITY CONTROL DEPARTMENT:
Allocation rate based on capcaity = $600000/60000DLH = $            10.00 per DLH
Allocation rate based on budgeted activity = $600000/40000MH = $            15.00 per DLH
3) UNUSED CAPACITY COST:
Cut and polish department = $8*(25000-20000) = $         40,000
Quality control department = $10*(60000-40000) = $      2,00,000
4) Actual plant wide rate based on labor cost = ($210000+$590000)/$150000 = 533.33% of labor cost
Actual rate for cut and polish department = $210000/$21000 = $            10.00 per MH
Actual rate for quality control department = 590000/48000 = $            12.29 per DLH